### v2

@@ 4.1 Experimental Settings and Details
Tasks. We evaluated models in three tasks Xu et al. (2017): 1) Predicate Classification (PredCls): Predicting the predicate category given all ground-truth entity bboxes and categories. 2) Scene Graph Classification (SGCls): Predicting categories of the predicate and entity given all ground-truth entity bboxes. 3) Scene Graph Generation (SGGen): Detecting all entities and their pairwise predicates.
Metrics. We evaluated SGG models on three metrics: 1) Recall@K (R@K): It indicates the proportion of ground-truths that appear among the top-$K$ confident predicted relation triplets. 2) mean Recall@K (mR@K): It is the average of R@K scores which are calculated for each predicate category separately. 3) Mean: It is the average of all R@K and mR@K scores. Since R@K favors head predicates while mR@K favors tail predicates, the Mean can better reflect the overall performance of all predicates Li et al. (2022a).
Datasets. We conducted all experiments on two datasets: VG and GQA. 1) VG: We followed the widely-used strategy Xu et al. (2017) to split dataset (i.e., 75K/32K images for train/test, 150 object categories and 50 predicate categories). Besides, 5K images within the training set are sampled as val set following Tang et al. (2020). All predicate categories are divided into three groups {head, body, tail} based on the number of samples as Liu et al. (2019). 2) GQA: We followed the prior work Dong et al. (2022) to split dataset (i.e., 70%/30% of the images for train/test, 200 object categories and 100 predicate categories). Similarly, 5K images within the training set are sampled as val set. We utilized the same group category ratio as VG to divide all predicate categories into {head, body, tail} groups.
Implementation Details. Following the training protocol in prior SGG works Tang et al. (2020), we adopted the object detector Faster R-CNN with the ResNeXt-101-FPN Lin et al. (2017) backbone trained by Tang et al. (2020) to detect all the bounding boxes and extract their visual features. The parameters of the backbone were kept frozen during the training. This detector could achieve 28.14 mAP on the VG test set (i.e., using 0.5 IoU threshold for evaluation). To avoid messy hyperparamter tuning, most of the hyperparamters follow preivous works. More specifically, the hyperparameters $\lambda$ and $\gamma$ were set to 0.07 and 0.7 in MCFA (cf., Sec. 3.2). The weights of pattern, context and semantic similarity were 1.0, 1.0 and 0.01 in intrinsic-CFA (cf., Sec. 3.2.1). The hyperparameter $\alpha$ used to control the value range of adaptive mixup coefficient $\theta$ was set to 0.5 in extrinsic-CFA (cf., Sec. 3.2.2) The threshold $\sigma$ was set to 0.5 in spatial restriction (cf., Sec. 3.2.2). The $\beta$ was set to 0.1 to regulate the loss during training (cf., Sec. 3.4). In this paper, SGD optimizer was used to train the model. The batch size was set to 12 and the initial learning rate was set to 0.01. After the performance on the val set reached the plateau period, the learning rate would be decayed by 10 for two times. All experiments were carried out with PyTorch and NVIDIA 2080Ti GPU.
@@ 4.2 Comparison with State-of-the-Arts
Setting. Due to the model-agnostic nature, we equipped our MCFA with three strong two-stage SGG baselines: Motifs Zellers et al. (2018), VCTree Tang et al. (2019) and Transformer Vaswani et al. (2017), and they are denoted as Motifs+MCFA, VCTree+MCFA, and Transformer+MCFA, respectively. In addition, due to the limited diversity of tail predicate components Li et al. (2022d), it has a high correlation with the category of subject & object. Thus, we further equip our three models with the component prior knowledge collected from the dataset to further improve the performance of the tail predicates. And they are denoted as Motifs+MCFA, VCTree+MCFA, and Transformer+MCFA (More details about the priors are discussed in appendix).
@@
Baselines. We compared our methods with the state-of-the-art (SOTA) models in the VG dataset (Table 1 and Table 2) and the GQA dataset (Table 3). Specifically, these models can be divided into three groups: 1) Model-specific designs: Motifs, VCTree, Transformer, and BGNN Li et al. (2021b). 2) Model-agnostic trade-off methods: consider the performance of all predicates comprehensively (i.e., higher Mean), e.g., PCPL Yan et al. (2020), DLFE Chiou et al. (2021), BP-LSA Guo et al. (2021), NICE Li et al. (2022a), IETrans Zhang et al. (2022), CFA Li et al. (2023a), and DPL Jeon et al. (2024). 3) Model-agnostic tail-focused methods: improve tail predicates performance at the expense of excessively sacrificing the head (i.e., higher mR@K and R@50 less than 50.0% on PredCls), e.g., TDE Tang et al. (2020), CogTree Yu et al. (2021), RTPB Chen et al. (2022), PPDL Li et al. (2022c), GCL Dong et al. (2022), and HML Deng et al. (2022). For a fair comparison, we compared with the methods in the last two groups.
Figure 7: The examples of reasonable entities for the query triplet in intrinsic-CFA. The lefts are original entities of the query triplets. The rights are some alternate entity categories (red) and their samples (gray boxes) for each query triplet to replace.
Figure 8: The examples of generated triplets for image-level augmentation.
Quantitative Results on VG. Analysis of the trade-off methods presented in Table 1 highlights two key observations: 1) Compared to three strong baselines (i.e., Motifs, VCTree, and Transformer), MCFA significantly enhances performance on the mR@K metric across all three experimental settings. 2) MCFA strikes the most effective trade-off between R@K and mR@K, evidenced by the highest Mean score. Notably, it surpasses the previous SOTA trade-off method, CFA Li et al. (2023a), which itself adeptly minimized performance decline on head predicates (R@K) while preserving tail performance (mR@K), in this crucial Mean metric under all settings. Turning to the tail-focused methods detailed in Table 2, MCFA, which incorporates further strategies dedicated to improving tail predicate performance, achieves the highest mR@K values. It outperforms prominent SOTA tail-focused approaches, i.e., HML Deng et al. (2022) and CFA Li et al. (2023a), on the mR@K metric across all evaluated settings. Additional experimental analyses are provided in the appendix.
@@
Quantitative Results on GQA. As demonstrated in Table 3, MCFA also markedly enhances mR@K performance over three strong baselines (i.e., Motif, VCTree, and Transformer) on the large-scale GQA dataset. These results certify the generalizability and effectiveness of our method.
@@ 4.3 Ablation Studies
Component SGCls
IN EX-fg EX-bg mR@50 / 100 R@50 / 100 Mean
8.7 / 9.3 39.0 / 39.7 24.2
11.6 / 12.4 37.4 / 38.6 25.0
15.6 / 16.5 34.9 / 36.2 25.8
14.3 / 15.7 35.9 / 37.0 25.7
17.0 / 18.0 35.8 / 36.9 26.9
17.1 / 18.3 35.4 / 37.0 27.0
17.8 / 18.8 34.4 / 35.6 26.7
17.9 / 19.4 35.5 / 36.7 27.4
Table 4: Ablation study on each component in feature-level augmentation Krishna et al. (2017). Motifs Zellers et al. (2018) is used in all ablations on VG Krishna et al. (2017). IN: Replace the intrinsic features. Ex-fg: Mix up the extrinsic features of foreground triplets. Ex-bg: Mix up the extrinsic features of background triplets.
Similarity SGCls
@@
16.1 / 17.3 34.4 / 35.8 25.9
17.1 / 18.3 36.1 / 37.3 27.2
18.1 / 19.1 33.5 / 34.8 26.4
16.9 / 18.2 35.0 / 36.3 26.6
17.4 / 18.5 35.7 / 36.8 27.1
16.5 / 17.6 37.0 / 38.2 27.3
17.9 / 19.4 35.5 / 36.7 27.4
Table 5: Ablation study on each similarity in clustering of intrinsic-MCFA.
Feature-Level Augmentation. We assessed the impact of each component within our feature-level CFA using the Motifs Zellers et al. (2018) baseline under the SGCls setting. Feature-level CFA comprises three primary components: replacing intrinsic features (IN), mixing extrinsic features of foreground triplets (EX-fg), and mixing extrinsic features of background triplets (EX-bg). The results, reported in Table 8, lead to the following observations: 1) Employing only the IN component yields a marginal improvement in mR@K (e.g., 2.9%–3.1% gains) while slightly degrading R@K performance (a 1.1%–1.6% loss). This is likely attributable to the limited diversity among tail predicate triplets, meaning that feature variety remains constrained even after intrinsic feature replacement. 2) Both EX-fg and EX-bg components substantially enhance mR@K and also achieve R@K scores relative to the baseline. For instance, mR@100 increases to 18.0% from the baseline’s 9.3%, and R@100 maintains 18.8% with these components. 3) Combining all three components (IN, EX-fg, and EX-bg) achieves the optimal trade-off, i.e., the highest Mean score.
Similarity of Clustering in Intrinsic-CFA. We investigated the impact of three distinct similarity measures (pattern, context, and semantic) on clustering within intrinsic-CFA, utilizing the Motifs Zellers et al. (2018) model as a baseline under the SGCls setting. Results presented in Table 5 demonstrate that relying solely on pattern similarity yields the poorest SGG performance. This deficiency occurs since pattern-based similarity alone disregards crucial contextual and semantic information, which is essential for guiding meaningful cluster formation. For instance, while both a boat and a car may share a “ridable” pattern, substituting a car for a boat would be semantically invalid in a maritime context (e.g., on the sea), as a car cannot operate there. Conversely, when contextual and semantic similarities are integrated into the analysis, SGG performance becomes substantially more robust, culminating in the highest Mean.
$K_{p}$ SGCls
@@
10 17.5 / 18.9 35.3 / 37.1 27.2
20 17.9 / 19.4 35.5 / 36.7 27.4
50 17.3 / 18.4 35.5 / 37.3 27.1
Table 6: Ablation study on the number of predicate clusters $K_{p}$ of intrinsic-CFA.
Different Cluster Number. We performed $K_{p}$in {10, 20, 50} to evaluate the impact of the number of predicate clusters under SGCls setting with Motifs Zellers et al. (2018). As illustrated in Li et al. (2023a), since when entity cluster number $K_{e}$ set to 15, there is a larger selection range of replaceable entity categories and richer feature diversity of tail predicate triplets, we use same entity cluster number setting. All results are reported in Table 6. We find that $K_{p}$ = 20 works best in mR@K. The reason may be that when $K_{p}$ is smaller, such as 10, the predicate clusters might be too coarse, grouping semantically distinct predicates together. This over-generalization can lead to less precise entity selection by potentially merging entities that play truly different roles across these distinct predicates, thereby reducing the discriminativeness for specific relations. When $K_{p}$ is larger (e.g., 50), the predicate clusters may become overly granular, which could prevent the model from capturing broader semantic similarities between closely related predicates. Consequently, this limits the effectiveness of relation-conditioned entity merging as each predicate cluster becomes too specific and offers fewer possibilities for generalization.
$\theta$ SGCls
@@
0.0 16.9 / 17.9 35.2 / 36.5 26.6
0.5 17.1 / 18.6 36.0 / 37.2 27.2
1.0 11.6 / 12.4 37.4 / 38.6 25.0
Adap 17.9 / 19.4 35.5 / 36.7 27.4
Table 7: Ablation study on mixup parameter $\theta$ of extrinsic CFA. Motifs Zellers et al. (2018) is used in all ablation on VG Krishna et al. (2017).
@@
Mixup Parameter $\theta$ in Extrinsic-CFA. As mentioned in Sec. 3.2.2, $\theta$ indicates the proportion of selected features in final augmented features. We investigated $\theta$in {0.0, 0.5, 1.0, adap} under the SGCls setting with Motifs Zellers et al. (2018) in Table 7. When $\theta$ is too small, the query triplet has too much impact on the other triplets in the image of the context triplet, and when $\theta$ is too large, feature augmentation of the query triplet is not strong enough. Our adaptive mixup operation alleviates this by dynamically calibrating $\theta$ based on the semantic similarity between the query and context predicates. Such calibration facilitates the targeted enrichment of tail predicates using relevant contexts, while concurrently preventing feature distortion from less pertinent ones, ultimately enhancing robustness.
Effectiveness of Each Enhanced Strategy in MCFA. We verified the effectiveness of each enhanced strategy on the original CFA, as presented in Table 8. Key observations are as follows: 1) Employing relation-conditioned entity clustering effectively identifies more suitable entities, leading to an improvement in Recall while maintaining a comparable mean Recall. 2) The adaptive mixup operation enables tail instances to better leverage contextual information from head instances, thereby significantly boosting mean Recall. 3) The adoption of image-level augmentation can simultaneously enhance recall and mean recall, which proves the rationality of generated triplets. 4) Combining all strategies allows MCFA to achieve the best Mean performance.
Component SGCls
En-IN En-EX IMG mR@50 / 100 R@50 / 100 Mean
17.0 / 18.4 34.9 / 36.1 26.6
16.4 / 17.6 37.1 / 38.2 27.3
18.3 / 20.1 34.1 / 35.4 27.0
18.1 / 18.9 35.0 / 36.2 27.1
17.9 / 19.4 35.5 / 36.7 27.4
Table 8: Ablation study on each enhanced strategy in MCFA on the VG dataset Krishna et al. (2017). En-IN: Utilize relation-conditioned entity clustering in intrinsic-CFA. En-EX: Employ the adaptive feature augmentation in extrinsic-CFA. IMG: Use the image-level augmentation.
5 Qualitative Analysis
@@
Visualization of Reasonable Entities in Intrinsic-CFA. Fig. 7 shows some alternative entities for query triplets. Although the visual features of these candidate entities can differ significantly from the original, they remain reasonable substitutions within the query triplet. For example, while zebra samples may vary considerably in color and texture from a giraffe, both are suitable because they share the common behavior of “eating plants”. Replacing a giraffe with a zebra in such instances introduces new visual information, thereby enriching the feature diversity associated with the eating predicate.
Visualization of Generated Triplets for Image-level Augmentation. Fig. 8 displays some generated triplets (cf., Sec. 3.3), illustrating our method’s capability to generate diverse examples (e.g., handle-on back of-toilet) and to utilize the original query triplet’s bounding box for their supervision.
6 Conclusion
@@
In this paper, we revealed the drawbacks of existing re-balancing methods and discovered that the key challenge for unbiased SGG is to learn proper decision boundaries under the severe long-tailed predicate distribution. To address this, we proposed a model-agnostic MCFA framework to enrich the feature space of tail categories. It augments both intrinsic and extrinsic features of relation triplets from the feature-level and image-level. At the feature-level, a relation-conditioned entity clustering strategy was designed for intrinsic augmentation and an adaptive mix-up strategy for extrinsic augmentation, dynamically adjusting based on predicate similarity. For image-level augmentation, we devised a generative method leveraging powerful models like GLIGEN to create new visual appearances for triplets, thereby overcoming in-domain data limitations. Comprehensive experiments on the challenging VG and GQA datasets consistently demonstrated that MCFA’s effectiveness for unbiased SGG.
Data Availability All experiments are conducted on publicly available datasets; see the references cited.
Appendix
This appendix is organized as follows:
  • The hierarchical clustering process and results mentioned in Sec. 3.2.1 are shown in Sec. A.
  • A more detailed discussion of prior knowledge mentioned in Sec. 4.2 is shown in Sec. B.
  • More experimental comparisons and analyses are shown in Sec. C
Input: Predicate category set $\mathcal{C}_{P}=\{c_{p_{i}}\mid i=1,2,\ldots,N_{p}\}$;
Entity category set $\mathcal{C}_{E}=\{c_{e_{i}}\mid i=1,2,\ldots,N_{e}\}$;
Predicate cluster similarity function $Sim_{P}$;
Entity cluster similarity function $Sim_{E}$;
Target number of predicate clusters $K_{p}$;
Target number of entity clusters $K_{e}$ within each predicate cluster;
Output: Set of all predicate clusters $\mathcal{S}^{p}=\{s^{p}_{i}|i=1,2,\cdots,K_{p}\}$
Set of all entity clusters $\mathcal{S}^{e}=\{s^{e}_{i}|i=1,2,\cdots,K_{e}\}$ within each predicate cluster $s^{p}_{i}$
/* --- Step 1: Predicate Clustering --- */
/* Initialize each predicate category as a cluster */
for $i=1,2,\cdots,N_{p}$ do
       $s_{i}^{p}=\{c_{p_{i}}\}$
/* Initialize the size of each cluster $l^{p}_{i}$ and the similarity matrix $A^{p}_{sim}$/
for $i=1,2,\cdots,N_{p}$ do
       $l^{p}_{i}=1$
      for $j=1,2,\cdots,N_{p}$ do
             $A_{sim}^{p}(i,j)=Sim_{P}(s^{p}_{i},s^{p}_{j})$
            $A_{sim}^{p}(j,i)=A^{p}_{sim}(i,j)$
/* Merge the two most similar predicate cluster until the number of clusters is smaller than $K_{p}$/
while $\text{LEN}(\mathcal{S}^{p})>K_{p}$ do
       $s^{p}_{i},s^{p}_{j}=\text{SELECT\_MAX}(A^{p}_{sim}(i,j)/(l^{p}_{i}+l^{p}_{j}))$
      $s^{p}_{i}=\text{MERGE}(s^{p}_{i},s^{p}_{j})$
      $\mathcal{S}^{p}=\mathcal{S}^{p}-{s^{p}_{j}}$
      $l^{p}_{i}=l^{p}_{i}+l^{p}_{j}+1$
      $\text{UPDATE}(A^{p}_{sim})$
/* --- Step 2: Predicate Merging --- */
for each predicate cluster $s^{p}_{i}\in\mathcal{S}^{p}$ do
       for $c_{e_{i}}\in C_{E}$ do
             if $c_{e_{i}}$ is in the same triplet category with $c_{p_{i}}\in s^{p}_{i}$ then
                   $C^{s^{p}_{i}}_{E}\leftarrow c_{e_{i}}$
/* --- Step 3: Entity Clustering within each Predicate Cluster --- */
for each predicate cluster $s^{p}_{i}\in\mathcal{S}^{p}$ do
       /* Initialize each entity category as a cluster */
      for $i=1,2,\cdots,|C^{s^{p}_{i}}_{E}|$ do
             $s_{i}^{e}=\{c_{e_{i}}\}$
      /* Initialize the size of each cluster $l^{e}_{i}$ and the similarity matrix $A^{e}_{sim}$/
      for $i=1,2,\cdots,N_{e}$ do
             $l^{e}_{i}=1$
            for $j=1,2,\cdots,N_{e}$ do
                   $A_{sim}^{e}(i,j)=Sim_{E}(s^{e}_{i},s^{e}_{j})$
                  $A_{sim}^{e}(j,i)=A^{e}_{sim}(i,j)$
      /* Merge the two most similar entity cluster until the number of clusters is smaller than $K_{e}$/
      while $\text{LEN}(\mathcal{S}^{e})>K_{e}$ do
             $s^{e}_{i},s^{e}_{j}=\text{SELECT\_MAX}(A^{e}_{sim}(i,j)/(l^{e}_{i}+l^{e}_{j}))$
            $s^{e}_{i}=\text{MERGE}(s^{e}_{i},s^{e}_{j})$
            $\mathcal{S}^{e}=\mathcal{S}^{e}-{s^{e}_{j}}$
            $l^{e}_{i}=l^{e}_{i}+l^{e}_{j}+1$
            $\text{UPDATE}(A^{e}_{sim})$
return $\mathcal{S}^{e}$ for each predicate cluster $s^{p}_{i}$;
Algorithm 1 Relation-Conditioned Entity Clustering
Figure 9: The results of relation clustering on VG Krishna et al. (2017).
Figure 10: The results of entity clustering on VG Krishna et al. (2017).
Appendix A Relation-Conditioned Entity Clustering
As mentioned in Sec. 3.2.1, we use relation-conditioned entity clustering to mine the replaceable entity categories. The specific implementation is detailed in Algorithm 1, which involves two main hierarchical clustering stages: first for predicates, and subsequently for entities, conditioned on these predicate clusters. Both stages utilize a similar iterative merging procedure.
The cluster similarity functions, $Sim_{P}$ for predicates and $Sim_{E}$ for entities, are the weighted sum of pattern, context, and semantic similarities. In practice, we treat clusters as a collection of multiple entity categories, and the pattern and context similarity can be transferred directly from class-level to cluster-level. For semantic similarity, we take the mean similarity for the categories between clusters.
After initializing the variables, the predicate and entity clustering stages share a similar logic in each iteration. First, we select the two most similar clusters (SELECT_MAX) based on similarity matrix ($A^{p}_{sim}$ for predicates or $A^{p}_{sim}$ for entities) and size of cluster ($l^{p}$ or $l^{e}$). Given that the distribution of different relation triplets in datasets is often highly imbalanced, we follow Wei et al. (2020) to employ the cluster size as a penalty term. This aims to prevent categories (whether predicate or entity) with a large proportion of samples from unduly dominating the clustering process. Subsequently, we merge the predicate or entity categories in cluster $s^{p}_{j}$ or $s^{e}_{j}$ into $s^{p}_{i}$ or $s^{e}_{i}$ (MERGE), remove cluster $s^{p}_{j}$ / $s^{e}_{j}$, and update the corresponding cluster size. Since the merged cluster has changed, we recalculate the similarity of other clusters to it and then update $A^{p}_{sim}$ / $A^{e}_{sim}$ (UPDATE). The above steps are repeated until the number of clusters is equal to $K^{p}$ for predicates or $K^{e}$ for entities.
The predicate clustering result and entity clustering results are shown in Fig. 9 and Fig. 10, respectively. Obviously, some predicate and entity categories that share common pattern are grouped into the same cluster, e.g., lying on and laying on (predicates), and dog and cat (entities, grouped within relevant predicate cluster contexts).
Figure 11: Performance(%) comparison between Motifs Zellers et al. (2018) and Motifs+CFA over all predicates on test set of VG Krishna et al. (2017). The orange area denotes the predicate distribution of training set.
Appendix B Component Prior Knowledge
As discussed in prior SGG works Li et al. (2022d), the statistic prior of the predicate distribution under a given condition can improve the mR@K performance of the unbiased SGG. In the inference phase, we calculate statistic component prior $b_{s,o,r}$, and add it to the predicted logits to predict predicate category:
${{{b}}_{s,o,r}}=-\log\frac{{{count_{s,o,r}}}}{{\sum\nolimits_{i=1}^{H}{{count_{s,o,i}}}}},$ (19)
where $H$ is the number of predicate categories. As for the prior “Triplet”, “Subject”, and “Object”, $count_{s,o,r}$ is the amount of the triplets whose predicate category is $r$ in the training set given subject-object pair, subject, and object respectively.
Appendix C Extra Comparisons and Analyses
Strategy PredCls
mR@50 / 100 R@50 / 100 Mean
Motifs Zellers et al. (2018) 16.5 / 17.8 65.6 / 67.2 41.8
+Reweight Tang et al. (2020) 30.8 ↑14.3 / 34.5 ↑16.7 36.1 / 40.4 35.5
+Resample Tang et al. (2020) 18.5 ↑2.0 / 20.0 ↑2.2 64.6 / 66.7 42.5
+MCFA 35.8 ${}_{{\color[rgb]{0.1725,0.7059,0.3921}\definecolor[named]{pgfstrokecolor}{rgb}{0.1725,0.7059,0.3921}\uparrow\textbf{18.7}}}$ / 38.3 ${}_{{\color[rgb]{0.1725,0.7059,0.3921}\definecolor[named]{pgfstrokecolor}{rgb}{0.1725,0.7059,0.3921}\uparrow\textbf{20.5}}}$ 54.3 / 56.8 46.3
Table 9: Performance (%) of re-balancing strategies and MCFA on VG Krishna et al. (2017).
Figure 12: The results of scene graphs generated by Motifs (blue) and Motifs-MCFA (green) on VG Krishna et al. (2017). Green predicates are correct (i.e., match GT), brown predicates are acceptable (i.e., does not match GT but still reasonable), and purple predicates are more informative (i.e., does not match GT and more reasonable).
Figure 13: The t-SNE visualization of randomly sampled instances of two predicates on the feature space before and after MCFA.
C.1 Comparison with Re-balancing Methods.
To demonstrate the superiority of MCFA compared with the prevalent re-balancing methods (i.e., reweight and resample) Tang et al. (2020), we conducted the three strategies on the baseline Motifs Zellers et al. (2018). The results under the PredCls setting are reported in Table 9. From the results, we can observe: 1) The Motifs baseline can achieve the best R@K. However, the high R@K is mainly due to the frequency bias of the dataset Tang et al. (2020), and they suffer severe drops in tail predicates. 2) The reweight method achieves better performance at mR@K, but it also sacrifices the performance of head predicates excessively, resulting in a low Mean. 3) The resample method keeps R@K high, but the improvement of mR@K is slight. The reason is that those decision boundaries may still be biased toward the head. 4) MCFA achieves the highest mR@K and maintains high R@K, i.e., it achieves the best trade-off over different predicate categories (highest performance on Mean).
C.2 Comparison over All Predicates
To further demonstrate the performance of each predicate, we displayed the predicate distribution of the training set in the VG dataset Krishna et al. (2017) and the performance of Motifs Zellers et al. (2018) and Motifs+MCFA on R@100 for each predicate category in Fig. 11. Obviously, our approach slightly compromises the performance of the head predicates (e.g., wearing), but greatly improves the tail predicates (e.g., flying in). This proves the superiority of our method in considering the performance of all predicates.
C.3 Qualitative Results.
Figure 12 shows some qualitative results generated by Motifs Zellers et al. (2018) and Motifs+MCFA under PredCls setting. We can observe that MCFA can not only predict more accurate predicates (e.g., on vsparked on and on vs.walking on), but also more fine-grained and informative predicates (e.g., holding vsusing).
C.4 Quality of Compositional Features.
We visualized the t-SNE distributions of features (from the previous layer of classification) in Fig. 13, demonstrating that our MCFA enhances the distinctiveness of tail predicates in feature space, i.e., it aids in finding suitable decision boundaries and does not harm to training, resulting in improved performance.

### v3

@@ 4.1 Experimental Settings and Details
Figure 6: The examples of reasonable entities for the query triplet in intrinsic CFA. The lefts are original entities of the query triplets. The rights are some alternate entity categories (red) and their samples (gray boxes) for each query triplet to replace.
Component PredCls
IN EX-fg EX-bg mR@50 / 100 R@50 / 100 Mean
16.5 / 17.8 65.6 / 67.2 41.8
19.3 / 21.2 64.7 / 66.6 43.0
25.6 / 27.8 63.0 / 64.8 45.3
23.9 / 26.3 63.3 / 65.7 44.8
27.2 / 29.3 61.9 / 64.3 45.7
27.5 / 30.0 61.7 / 64.1 45.8
27.8 / 30.3 60.7 / 63.4 45.6
35.7 / 38.2 54.1 / 56.6 46.2
Figure 7: Ablation study on each component on VG [20]. IN: Replace the intrinsic features. Ex-fg: Mix up the extrinsic features of foreground triplets. Ex-bg: Mix up the extrinsic features of background triplets.
Tasks. We evaluated models in three tasks [40]: 1) Predicate Classification (PredCls): Predicting the predicate category given all ground-truth entity bboxes and categories. 2) Scene Graph Classification (SGCls): Predicting categories of the predicate and entity given all ground-truth entity bboxes. 3) Scene Graph Generation (SGGen): Detecting all entities and their pairwise predicates.
Metrics. We evaluated SGG models on three metrics: 1) Recall@K (R@K): It indicates the proportion of ground-truths that appear among the top-$K$ confident predicted relation triplets. 2) mean Recall@K (mR@K): It is the average of R@K scores which are calculated for each predicate category separately. 3) Mean: It is the average of all R@K and mR@K scores. Since R@K favors head predicates while mR@K favors tail predicates, the Mean can better reflect the overall performance of all predicates [21].
Datasets and Implementation Details. We conducted all experiments on two datasets: VG [20] and GQA [16]. More details of datasets and implementation are in the appendix.
@@ 4.2 Comparison with State-of-the-Arts
Setting. Due to the model-agnostic nature, we equipped our CFA with three strong two-stage SGG baselines: Motifs [44], VCTree [36] and Transformer [37], and they are denoted as Motifs+CFA, VCTree+CFA, and Transformer+CFA, respectively. In addition, due to the limited diversity of tail predicate components [28], it has a high correlation with the category of subject & object. Thus, we further equip our three models with the component prior knowledge collected from the dataset to further improve the performance of the tail predicates. And they are denoted as Motifs+CFA, VCTree+CFA, and Transformer+CFA (More details about the priors are discussed in appendix).
@@
Baselines. We compared our methods with the SOTA models in the VG dataset (Table 1, Table 2) and the GQA dataset (Table 3). Specifically, these models can be divided into three groups: 1) Model-specific designs: Motifs, VCTree, Transformer, and BGNN [26]. 2) Model-agnostic trade-off methods: consider the performance of all predicates comprehensively (i.e., higher Mean), e.g., PCPL [41], DLFE [5], BP-LSA [11], NICE [21], and IETrans [45]. 3) Model-agnostic tail-focused methods: improve tail predicates performance at the expense of excessively sacrificing the head (i.e., higher mR@K and R@50 less than 50.0% on PredCls), e.g., TDE [35], CogTree [42], RTPB [1], PPDL [27], GCL [10], and HML [7]. For a fair comparison, we compared with the methods in the last two groups.
Quantitative Results on VG. From the results of trade-off methods in Table 1, we can observe that: 1) Compared to the three strong baselines (i.e., Motifs, VCTree and Transformer), CFA can significantly improve model performance on mR@K metric over all three settings. 2) CFA can achieve the best trade-off between R@K and mR@K, i.e., highest Mean, and surpass the SOTA trade-off method NICE [21] in Mean metric under all settings. CFA shows minimal performance degradation on the head predicates (c.f. R@K) while maintaining the performance of the tail (c.f. mR@K), demonstrating the superiority of CFA considering all predicates. From the results of tail-focused methods in Table 2, we can observe that: after further implementing strategies to improve tail performance, CFA can achieve the highest mR@K and exceed the SOTA tail-focused method HML [7] on mR@K metric under all settings. More experiment analysis is in the appendix.
@@
Quantitative Results on GQA. From the results of Table 3, we can observe that: CFA can also greatly improve the mR@K of three strong baselines (i.e., Motif, VCTree and Transformer) on the large dataset GQA, which proves the universality and effectiveness of our method.
@@ 4.3 Ablation Studies
Similarity PredCls
@@
33.4 / 36.2 52.0 / 54.0 43.9
34.6 / 37.3 55.2 / 57.0 46.0
35.3 / 38.0 51.4 / 54.1 44.7
35.3 / 37.8 52.6 / 55.2 45.2
35.2 / 37.9 54.0 / 56.2 45.8
34.5 / 36.9 55.2 / 57.5 46.0
35.7 / 38.2 54.1 / 56.6 46.2
$K$ PredCls
@@
15 35.7 / 38.2 54.1 / 56.6 46.2
40 33.5 / 36.0 55.3 / 57.8 45.7
150 33.2 / 35.7 55.7 / 57.9 45.6
$\theta$ PredCls
@@
0.0 27.1 / 29.5 62.2 / 64.3 45.8
0.3 28.5 / 31.0 60.8 / 63.3 45.9
0.5 35.7 / 38.2 54.1 / 56.6 46.2
0.7 32.9 / 35.2 50.4 / 53.0 42.9
1.0 19.3 / 21.2 64.7 / 66.6 43.0
Table 4: Ablation studies on the different hyperparameters of each component of CFA. Motifs [44] is used in all ablation on VG [20].
Effectiveness of Each Component. We evaluated the importance of each component of CFA based on Motifs [44] under the PredCls setting. There are three components of CFA: replace the intrinsic features (IN), mix up the extrinsic features of the foreground triplets (EX-fg), and mix up the extrinsic features of the background triplets (EX-bg). As reported in Table 7, we have the following observations: 1) Using only IN can slightly improve mR@K (e.g., 2.8% $\sim$ 3.4% gains) and slightly hurts R@K (0.6% $\sim$ % 0.9 loss). The reason is that there are not enough tail predicate triplets, resulting in the feature diversity still being limited after replacing the intrinsic features. 2) Both Ex-fg and Ex-bg can significantly improve mR@K and keep competitive R@K compared to baseline (e.g., 17.8% vs. 30.3% in mR@100, and 67.2% vs. 63.4% in R@100). 3) Combining all components allows for the best trade-off, i.e., the highest Mean.
Similarity of Clustering in Intrinsic CFA. We analyzed the influence of three kinds of similarity (pattern, context, and semantic similarities) under the PredCls setting with baseline model Motifs [44]. From Table 4(a), we can observe that using only pattern similarity shows the worst performance, since it ignores the contextual and semantic information which gives crucial guidance for selection. For example, both boat and car can be ridden, but we cannot replace boat with car because car can’t appear in the sea. Once the other two similarities are fused, the SGG performance becomes robust and the highest Mean is achieved.
Different Clusters $K$ of Clustering in Intrinsic CFA. We performed $K$ $\in$ {15, 40, 150} to evaluate the impact of the number of clusters under PredCls setting with Motifs [44]. All results are reported in Table 4(b). We find that KK=15 works best in mR@K. The reason may be that when KK=15, there is a larger selection range of replaceable entity categories and richer feature diversity of tail predicate triplets.
@@
Mixup Parameter $\theta$ in Extrinsic CFA. As mentioned in Sec. 3.2.2, $\theta$ indicates the proportion of selected features in final augmented features. We investigated $\theta$ $\in$ {0.0, 0.3, 0.5, 0.7, 1.0} under the PredCls setting with Motifs [44] in Table 4(c). When $\theta$ is too small, the query triplet has too much impact on the other triplets in the image of the context triplet, and when $\theta$ is too large, feature augmentation of the query triplet is not strong enough. To better trade-off the performance on different predicates, we set $\theta$ to $0.5$.
@@
Visualization of Reasonable Entities in Intrinsic CFA. Figure 7 shows some alternative entities for query triplets. The visual features of the candidate entities are quite different from the original entity, but are still reasonable for the query triplet, e.g., the samples of zebra differ greatly in color and texture from the giraffe, but they can all eat plants. Replacing the giraffe with the zebra can provide new features to enrich the feature diversity of eating.
5 Conclusion and Future Work
@@
In this paper, we revealed the drawbacks of existing re-balancing methods and discovered that the key challenge for unbiased SGG is to learn proper decision boundaries under the severe long-tailed predicate distribution. Thus, we proposed a model-agnostic CFA framework that can enrich the feature space of tail categories by augmenting both intrinsic and extrinsic features of the relation triplets. Comprehensive experiments on the challenging VG and GQA datasets showed that CFA significantly improves the performance of unbiased SGG. In the future, we would like to extend CFA to compose new triplets features by fusing features from open-vocabulary categories or images from other domains.
Acknowledgement. This work was supported by the National Key Research & Development Project of China (2021ZD0110700), the National Natural Science Foundation of China (U19B2043, 61976185), and the Fundamental Research Funds for the Central Universities(226-2023-00048). Long Chen was supported by HKUST Special Support for Young Faculty under Grant F0927.